OpenStudy (anonymous):

9) if 2x=sec theta and 2/x=tan theta , find 2(x^2=1/x^2)

myininaya (myininaya):

find what?

OpenStudy (oblivious):

if the equals is supposed to be a minus sign (which is the only way this problem makes much sense) \[x=\sec(\theta)/2\]and\[x=2/[\tan(\theta)]\] thus \[2*(x^2-1/x^2) = 2*(\sec^2(\theta)/4-\tan^2(\theta)/4) = 1/2*(1) = 1/2\]

OpenStudy (anonymous):

yea..sorry typing mistake!